Calculus I Fall 2001 MA113:13-15, 19-21 Russell Brown

• Homework#14, due Monday, 3 December 2001. §4.4 #18, 28, 72 82. §4.5 #40, 42, 44, 64. (Apparently, there were two homework #12’s. Thus we must skip #13, in order to get back on track.)

• Homework H. due Wednesday, 5 December 2001. §5.2 #48.

• Homework 15, not to be collected. §5.1 #9, 19, 25, 29. §5.2 #11, 33, 49, 51. §5.3 #9, 25, 33, 41.

December 2, 2001